OpenStudy (anonymous):

What is the surface area of the rectangular prism? 126 cm2 63 cm2 14 cm2 590 cm2

OpenStudy (anonymous):

rectangular prism

Directrix (directrix):

Surface area of a right prism is the perimeter of the base times the height PLUS 2* area of base.

Directrix (directrix):

SA = (5+5+3+3)*6 + 2* (5*3) SA = ?

OpenStudy (anonymous):

126

Directrix (directrix):

That is what I got.

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

What is the volume of the rectangular prism? 56 cm3 40 cm3 11 cm3 28 cm3

Directrix (directrix):

Volume of right prism is the area of the base times the height. V = (5*2) * 4 V =

OpenStudy (anonymous):

40

OpenStudy (anonymous):

thank you soo much :D

Directrix (directrix):

Okay, that's a wrap.